when is to use R=0.287 kj/kg.K and Cv=0.718?
when is to use gamma = 1.4
Answer #1

R=0.287kj/kgk and cv=.718kj/kg k used when working medium is ideal gas.Air act like ideal gas so we can use the constant R and cv Values that are mentioned above.Gamma(r) value calculated by r=(1+2/n) ,where n = number of degree of freedom of atom ,for diatomic gas Dof = 5. r=(1+2/5) r=1.44 .Air is act like diatomic gas so we can use r=1.4 for air.Examples of diatomic gases H2 ( two hydrogen atoms),O2 ( two oxygen atoms ) ,CO(1carbon + 1 oxygen),HCL( one hydrogen and one chlorine).